OpenStudy (anonymous):

The cost of a pack of batteries, after a 5% tax is $8.40 . What was the price before tax?

OpenStudy (anonymous):

x +0.05x = 8.4 = 1.05x

OpenStudy (anonymous):

Is there a proportion for it?

OpenStudy (anonymous):

x = 8.4/1.05 = 8$

OpenStudy (anonymous):

Wait, how'd you get that equation? @Euler271

OpenStudy (anonymous):

let x = original price original price + tax = $8.4 tax is calculated by a percentage of the value of the item (i know you know :P) so it's x + (5% * x) = 8.4 and 5% = 5/100 = 0.05 so it's x + 0.05x = 8.4